find the indicated outputs for f(x)=3x²-2x
f(0)=
f(-1)=
f(2)=

Answer by stanbon(75887)

find the indicated outputs for
(x)=3x²-2x
f(0)= 3*0^2-2*0 = 0-0 = 0
---------------------------------
f(-1)= 3*1^2-2*1 = 3-2 = 1
---------------------------------
f(2)= 3*2^2-2*2 = 12-4 = 8
===============================
Cheers,
stan H.